What is the mass, in grams, of 28.46 mL of acetone? Acetone (nail polish remover) has a density of 0.7857 g/cm3.
1
Expert's answer
2015-09-25T04:55:42-0400
r = m/V
m = V*r
V = 28.46 ml = 28.46 cm^3
m = 28.46*0.7857 = 22.36 g
Answer: 22.36 g